OpenStudy (hulagirlmb):

I need help understanding how to find the slope/y intercept of a linear equation

Nnesha (nnesha):

do you have any equation ??

OpenStudy (anonymous):

Linear equation?

OpenStudy (anonymous):

Like the equation y = mx + b?

OpenStudy (anonymous):

If you are given the equation y = mx + b, you can find the slope and y-intercept. Because m = slope, and b = y-intercept.

OpenStudy (hulagirlmb):

\[2y + 4 =3x\]

Nnesha (nnesha):

slope you have to use formula  to find slope

Nnesha (nnesha):

solve for y first

Nnesha (nnesha):

subtract 4 both side

OpenStudy (hulagirlmb):

I already did that

Nnesha (nnesha):

okay what you got after taht

OpenStudy (hulagirlmb):

It looks like this \[2y = 3x-4\]

Nnesha (nnesha):

yes now divide both side by 2

OpenStudy (anonymous):

Oh, now divide all terms by 2: \(y = \dfrac{3}{2}x - 2\) Now it's int he form of y = mx + b, where m is slope, and b is y-intercept. So what's the slope and y-intercept?

OpenStudy (hulagirlmb):

So does it look like this \[y= 3/2x-2\]

OpenStudy (hulagirlmb):

Oh lol ok. The slope is 3/2 and the y intercept is -2

Nnesha (nnesha):

yes right now slope intercept form equation is y = mx +b where m is slope which is always with x

Nnesha (nnesha):

yes right

OpenStudy (hulagirlmb):

Thank you

Nnesha (nnesha):

\(\huge\color{blue}{my ~ pleasure }\)